How many moles are in 9.6x10^16 molecules of phosphorus trichloride (Pcl3)
1
Expert's answer
2020-05-28T10:32:29-0400

"moles=\frac{number of molecules}{Avogadro constant}=\frac{9.6\times10^{16}}{6.02\times10^{23}}=1.6\times 10^{-7} mol"